### v1

# Prismatic-Bending Transformable (PBT) Joint for a Modular, Foldable Manipulator with Enhanced Reachability and Dexterity

Abstract
Robotic manipulators, traditionally designed with classical joint-link articulated structures, excel in industrial applications but face challenges in human-centered and general-purpose tasks requiring greater dexterity and adaptability. Addressing these limitations, we introduce the Prismatic-Bending Transformable (PBT) Joint, a novel design inspired by the scissors mechanism, enabling transformable kinematic chains. Each PBT joint module provides three degrees of freedom—bending, rotation, and elongation/contraction—allowing scalable and reconfigurable assemblies to form diverse kinematic configurations tailored to specific tasks. This innovative design surpasses conventional systems, delivering superior flexibility and performance across various applications. We present the design, modeling, and experimental validation of the PBT joint, demonstrating its integration into modular and foldable robotic arms. The PBT joint functions as a single SKU, enabling manipulators to be constructed entirely from standardized PBT joints without additional customized components. It also serves as a modular extension for existing systems, such as wrist modules, streamlining design, deployment, transportation, and maintenance. Three sizes—large, medium, and small—have been developed and integrated into robotic manipulators, highlighting their enhanced dexterity, reachability, and adaptability for manipulation tasks. This work represents a significant advancement in robotic design, offering scalable and efficient solutions for dynamic and unstructured environments.
Index Terms:
Transformable Robot, Robotic Joint, Robotic Manipulator, Modular Robot, Robotic Manipulation
I Introduction
Figure 1: (A) The modular manipulator composed of two PBT joints and a dexterous end-effector. (B) and (C) show the PBT wrist. The PBT wrist is a small-sized PBT joint with a suction cup at its tip. (B) is performing the revolute motion. (C) demonstrates the prismatic motion.
ROBOTIC manipulators are indispensable tools that empower robots to interact with the physical world, much like human arms facilitate our own interactions. Traditional robotic manipulators are built on two fundamental joint types—prismatic and revolute—connected by rigid links to form desired kinematic chains [1]. These systems have resulted in six widely recognized geometric configurations: articulated, spherical, SCARA, cylindrical, Cartesian, and parallel manipulators [2]. Among these, articulated manipulators are particularly prevalent due to their intuitive kinematics, making them well-suited for high-accuracy, high-payload tasks [3, 4]. However, human-centered environments, such as domestic or service applications, present unique challenges that prioritize adaptability, dexterity, obstacle avoidance, and reachability over traditional metrics such as payload capacity and precision [5, 6, 7, 8]. Additionally, while robots performing tasks in structured environments primarily focus on avoiding obstacles at the end-effector, in unstructured environments, the body configuration of the manipulator is significantly influenced by the complexity and dynamics of the environment. Achieving greater flexibility in body configuration while maintaining terminal manipulability is crucial for enabling robots to effectively perform tasks in the unstructured settings [2, 5, 9].
To enhance the dexterity and reachability of robotic manipulators in unstructured environments and human-centered tasks, researchers have explored two distinct solutions: continuum manipulators and foldable robots, both offering innovative alternatives to traditional articulated manipulators. Continuum manipulators, inspired by biological structures such as elephant trunks and octopus arms, achieve smooth, adaptive motion through continuously deformable materials or mechanisms [10, 11]. These systems provide multi-directional bending, enabling exceptional obstacle avoidance and adaptability in unstructured environments [12, 13]. Our previous work has also demonstrated the potential of continuum robots to enhance dexterity in confined spaces. For example, a soft wrist design utilizing tendon differential actuation enables human wrist-like dexterity and manipulability for teleoperated swab sampling [14, 15]. By incorporating a tendon-jamming mechanism, the stiffness of the continuum manipulator becomes adjustable while maintaining excellent motion dexterity for operations in confined spaces [16].
Foldable manipulators leverage the unique capabilities of foldable mechanisms to enhance the reachability and adaptability of robotic systems [17]. By allowing their structures to extend, fold, or reshape, these manipulators offer exceptional versatility and can be tailored to meet task-specific requirements. The principles of modularity and foldability enable quick deployment, efficient space-saving storage, and effective obstacle avoidance, making them ideal for use in dynamic and constrained environments [18]. Foldable manipulators have demonstrated their unique advantages in various applications. For instance, a foldable manipulator integrated with UAVs showcases its ability to extend into confined spaces for tasks such as object retrieval and inspections, highlighting its potential in challenging scenarios like search and rescue operations [19]. Furthermore, foldable manipulators can effectively merge locomotion and manipulation functions, making them a promising solution for space exploration. By combining these capabilities, such manipulators can navigate and interact with complex extraterrestrial environments, as demonstrated in ReachBot’s integration for extraterrestrial missions [20].
Although continuum and foldable manipulators each offer distinct advantages, neither approach successfully integrates multi-directional bending with foldability into a single system—a promising avenue for significantly enhancing the dexterity and reachability of robotic manipulators.
To address this need, this paper introduces the Prismatic-Bending Transformable (PBT) Joint, a novel mechanism that combines multi-directional bending flexibility with foldability inspired by the scissors mechanism. Each PBT Joint provides three degrees of freedom—bending, rotation, and elongation/contraction—with transformable prismatic and bending modes. This design enables the assembly of modular and foldable robotic manipulators that are highly adaptable to unstructured environments and diverse task requirements. Details of the design, modeling, and experimental validation of the Prismatic-Bending Transformable (PBT) Joint are discussed, showcasing its ability to integrate seamlessly into robotic arms while significantly enhancing their dexterity, manipulability, and obstacle avoidance. Through quantitative analysis of its dexterity and manipulability, the PBT Joint integrated manipulator demonstrates superior performance in adapting to diverse tasks and complex environments. Serving as a single standardized SKU, the PBT Joint enables the construction of complete manipulators without requiring additional customized components. It also functions as a modular extension for existing robotic systems, such as a PBT wrist module, further broadening its applicability. To meet various application requirements, three sizes of PBT Joints—large, medium, and small—have been developed and tested in different robotic configurations. These implementations underscore the joint’s adaptability, reconfigurability, and effectiveness, positioning it as a promising solution for general-purpose robotic tasks and dynamic, human-centered applications.
The contributions of this work are summarized in three parts:
  • Novel Joint Design: Development of the Prismatic-Bending Transformable (PBT) Joint, a modular unit integrating multi-directional bending, rotation, and elongation/contraction with a 3D direction maintenance mechanism, offering adaptable dexterity and task-specific customization.
  • Enhanced Robotic Manipulator Architecture: Development of a modular and foldable robotic manipulator architecture using PBT joints, enabling enhanced dexterity, manipulability, and obstacle avoidance, with applications as an extension wrist or a foldable two-joint manipulator for unstructured environments.
  • Detailed analysis, modeling, planning, and experimental validation demonstrating the feasibility, performance, and versatility of the PBT joint in enabling robotic arms to perform a wide range of tasks with enhanced dexterity and manipulability.
The paper is organized as follows: Section II introduces the transformation between prismatic and bending modes in the PBT design. Section III details the PBT Joint’s design and its integration into modular manipulators. Section IV analyzes the kinematics and manipulability of a dual-unit system. Section V presents experimental validations for single joints and modular systems, followed by the conclusion and future work.
Figure 2: Concept of the Prismatic-Bending transformable (PBT) joint. (A) The motion illustration of the PBT joint. The first three stages are linear motion, while the third and fourth stages are revolute motion. The desired joint angles in the moving process are shown by the coordinates $S_{1},S_{2},andS_{3}$. (B) Serial combination of two PBT joints. The ’P’ represents prismatic, while ’B’ represents bending.
II Concept of PBT Joint
This section introduces the motion and degrees of freedom (DOFs) of a single PBT joint and its modular extension. As illustrated in Figure 2A, the proposed PBT joint employs a scissors-inspired mechanism to achieve reliable prismatic and bending motion, providing two correlated DOFs [21, 22, 23, 24]. This mechanism forms the basic PBT joint, enabling prismatic and bending transformations. Self-rotation is achieved by integrating a perpendicular rotation actuator with the core scissors mechanism [25].
Specifically, assuming the base frame ($\text{S}_{1}$) is fixed, driving Links #1 and #2 near the base causes the distal endpoint ($\text{S}_{3}$) to perform linear motion within the range $(0,2L]$. When $\text{S}_{3}$ reaches the farthest point, the PBT structure transitions to the serial chain, which is similar to a traditional revolute joint. At this stage, applying torque to the $S_{2}$ enables rotational motion between the coincident Links #3, #4 and Links #1, #2.
For modular extension, Figure 2B illustrates two serially connected PBT joints, enabling four transformable kinematic modes: (i) prismatic-prismatic (PP) motion, (ii) prismatic-bending (PB) motion, (iii) bending-prismatic (BP) motion, and (iv) bending-bending (BB) motion. Each stage operates independently, allowing versatile motion configurations.
Figure 3: 3D Direction Maintenance Mechanism. (A) Illustrates the Direction Maintenance Challenge in the PBT Joint. Using an open-loop link connection to implement the PBT joint causes the direction change at the PBT joint’s end. (B) PBT Joint with 3D Direction Maintenance Mechanism: Each end of the base links is equipped with a direction maintenance structure. (C) 3D Direction Maintenance Structure: The mechanism effectively withstands overturning torque across all three axes.
III Design of PBT Joint and Modular Manipulator Architecture
In this section, we present the detailed design of the PBT joint and its resulting modular manipulator. This includes the direction-maintenance mechanism of the PBT joint, the design and actuation principles, and the assembly process of the modular manipulator.
Figure 4: Structure and Assembly of the Two-Module PBT Joint Modular Manipulator (A) Partial sectional front view of the modular manipulator: 1 represents the timing belt, 2 is the servo motor driving the linear motion of the PBT joint, and 3 represents the driving gear of the PBT joint directly connected to the servo motor. (B) Full view of the modular manipulator: 4 represents the servo motors driving the revolute motion of the PBT joint, 5 represents the servo motor driving the revolute motion of the medium PBT joint, and 6 represents the servo motor driving the linear motion of the medium PBT joint. (C) Partial sectional view of the modular manipulator: The simplified link and joint structure are illustrated for the large PBT joint. (D) Partial exploded sectional view of the modular manipulator: From bottom to top are the large revolute module, large PBT module, medium revolute module, and medium PBT module. (E) Two potential structures for folding are shown. Force singularity occurs in the structure in the second row.
III-A Direction Maintenance Mechanism
Although Figs. 2A and 2B intuitively illustrate the relative directions of the coordinates of the PBT joint, translating this concept into practical implementation introduces challenges related to direction maintenance. In traditional joint-link mechanisms, the
$link\rightarrow joint\rightarrow link$
architecture allows for seamless assembly in both serial and parallel robotic structures. However, in the PBT joint, multiple links are connected at a single joint, forming a
$\begin{bmatrix}link\\ link\\ \ldots\end{bmatrix}\begin{matrix}\searrow\\ \rightarrow\\ \nearrow\end{matrix}joint\begin{matrix}\nearrow\\ \rightarrow\\ \searrow\end{matrix}\begin{bmatrix}link\\ link\\ \ldots\end{bmatrix}$
architecture. If the motor stator is connected to one link and the rotor to another, it creates a direction maintenance issue, as shown in Fig. 3A. This setup disrupts the endpoint’s direction.
The 3D direction maintenance mechanism is critical for ensuring the stability and precise functionality of the PBT joint in three-dimensional applications. Without it, the connection and extension of PBT joints cannot maintain consistent directionality between modules, leaving the connections floating and rotatable, which severely compromises the robot’s maneuverability. This mechanism effectively addresses the direction maintenance challenge by mechanistically ensuring alignment and withstanding torque along the x, y, and z axes, delivering robust and customizable performance to enhance the joint’s reliability in unstructured and confined environments.
As illustrated in Figure 3B and 3C, the proposed 3D direction maintenance mechanism consists of driving gears, slave gears, base links, and arm links. The driving and slave gears are securely fixed to the base links, and by rotating one or both driving gears (#1 and #2), the arm links near the base link open or close, facilitating the linear motion of the PBT joint. The interlocking properties of the differential mechanism ensure high stability and adaptability, making it a reliable solution for dynamic and complex spatial tasks. This design significantly improves the PBT joint’s performance in unstructured environments.
III-B Design and Actuation of the PBT Joint
Figure 4A and 4B provide a comprehensive illustration of the PBT joint’s structure, including the 3D direction maintenance mechanism with a gear reduction ratio, synchronized servo motor pairs, singularity-free foldable links, and a revolute joint with reduction. These four components are essential for the stable operation of the PBT joint under load. Customized design parameters are summarized in Tab. I.
The 3D direction maintenance mechanism, in conjunction with synchronized servo motors, enables the linear motion of the PBT joint (Fig. 4B). Building on the previously introduced 3D direction maintenance mechanism, we adjust the gear reduction ratio by varying the radii of the driving gear and the slave gear. For example, if $R_{1}:R_{2}=1$, a 90-degree motor rotation results in a $2L$ linear movement of the PBT joint’s endpoint. Additionally, we utilize two driving gears to further enhance the load-carrying capability during linear motion. As the direction maintenance mechanism employs a differential gear structure, driving gears can be positioned in any of the other four directions. We select a pair of opposing gears as the driving gears to balance the center of gravity. The servo motors are directly connected to the driving gears through couplings. Notably, the embedded synchronization feature of these servo motors eliminates the need for extensive compliance design or synchronization algorithms, allowing straightforward synchronized control of the two motors.
The singularity-free foldable links are an optimized variation of the closed-loop four-bar linkage structure. This optimization aims to achieve a higher folding ratio without compromising the functionality of the PBT joint (Fig. 4E). Considering the 3D direction maintenance mechanism occupies the central space and prevents fully overlapping folding, we adopt a stacked link structure. Furthermore, to avoid force singularities at the lower positions, which could hinder linear motion, the foldable links are designed with links #1 and #2 as straight links, and links #3 and #4 as L-shaped links. This design achieves effective folding while avoiding force singularities.
The revolute joint in the PBT joint comprises an electromagnetic clutch, a synchronous belt reduction mechanism, and a pair of servo motors (Fig. 4A). During linear motion, the revolute joint acts as a passive joint, with the clutch disengaged to ensure no interference between the motors for linear and revolute motion. When the system transitions to rotational mode (at the farthest linear position), the clutch engages. The synchronous belt transmission provides three key advantages: (1) reducing the inertia of the PBT joint by shifting the main weight closer to the base, (2) introducing a reduction ratio to increase joint torque, and (3) leveraging the compliance of the belt to compensate for any synchronization imperfections in the servo motors.
TABLE I: CUSTOMIZED PARAMETERS OF THE MODULAR MANIPULATOR
$Parameters$ $Value$
Large PBT Joint
Reduction ratio of linear motion 3
Reduction ratio of revolute motion 2
Length of link 1 16 ($cm$)
Length of link 3 20 ($cm$)
Self-Weight 5 ($kg$)
Linear Payload 9 ($kg$)
Medium PBT Joint
Reduction ratio of linear motion 3
Reduction ratio of revolute motion 2
Length of link 1 10 ($cm$)
Length of link 3 12 ($cm$)
Self-Weight 2.2 ($kg$)
Linear Payload 4 ($kg$)
Small PBT Joint (PBT Wrist)
Reduction ratio of linear motion 2
Reduction ratio of revolute motion 1
Length of link 1 5 ($cm$)
Length of link 3 7 ($cm$)
Self-Weight 0.5 ($kg$)
Linear Payload 0.8 ($kg$)
III-C Assembly of Multi-PBT Joints for Modular Manipulator
The modular manipulator consists of multiple PBT joints and revolute joints, where revolute joints are independent rotary motors (Fig. 4B and Fig. 4C). The assembly of the modular manipulator allows for customization in two aspects: the sequence of joints and the size of the joints. These customizations are tailored based on task requirements and torque demands.
From the base to the end effector, the structure includes a revolute joint, a PBT joint, a revolute joint, and a PBT joint (Fig. 4D). Besides this configuration, other sequences may be more suitable for different tasks. For instance, a single PBT joint or a combination of a PBT joint and a revolute joint can serve as a dexterous wrist (Fig. 1B). A series of PBT joints can enable planar manipulator operations, while a custom configuration of PBT and revolute joints can create an extendable exoskeleton to potentially address current challenges in exoskeleton teleoperation. To accommodate varying torque demands, we provide three sizes of PBT joints: large, medium, and small. Their specific parameters are summarized in TableI. Taking the modular manipulator in this study as an example, the large and medium modules are used as the base and elbow, respectively, to meet the torque requirements of different positions (Fig. 4A).
IV Kinematics and Manipulability Analysis
This section establishes a comprehensive kinematics analysis for the modular manipulator. We first propose a detailed inverse kinematics (IK) solution for a two-unit modular unit arm, then showcase the extended reachability set of PBT joints with different modes.
IV-A Inverse Kinematics
Given the target end effector position $p_{t}$ and $M$ convex obstacles $\{\mathcal{O}_{m}\}_{m=1}^{M}$ in the task space, IK aims to find a collision-free joint configuration that reaches $p_{t}$ [26]. With two serially arranged modular units, there are four possible motion modes as shown in Fig. 2. Let $l_{1}$ and $l_{2}$ respectively be the arm link length of the lower and upper units. Use $t_{1}$ and $t_{2}$ to denote the thickness of the base revolute motors of the lower and upper units. $\phi_{1}$ and $\phi_{2}$ denotes the joint pose of the two base revolute joints, whereas $\theta_{1}$ and $\theta_{2}$ denotes the central revolute joint poses. When a unit is fully stretched, $\theta$’s are zero. For linear motion, the $\theta$ value should be non-negative.
Figure 5: Reachability and Manipulability Analysis (A) Reachability set of a single PBT joint: (i) and (ii) correspond to the prismatic and bending modes, respectively. (B) Single PBT joint (wrist) Mounted on a Collaborative Robot Arm: The reachability set in a compact environment is extended. (i) visualizes the environment, while (ii) and (iii) show workspaces with and without the mounted PBT joint. Scatter plots represent end-effector positions from 10,000 uniformly sampled joint configurations, excluding those with collisions. In (ii), sampling density is increased in the cavity area to better capture reachability. (C) Translational Manipulability Distribution in Task Space: (i)-(iv) correspond to the PP, PB, BP, and BB modes defined in Fig. 2. Translational manipulability quantifies the modular manipulator’s ability to move the end-effector in different directions in the task space, calculated using (5). (D) Workspace in a Compact Environment: (i) visualizes the environment. (ii) and (iii) show the BP and BB mode workspaces, represented by scatter plots from a top-down view. The scatter plots are generated from 30,000 randomly sampled joint configurations, excluding collisions. (E) Workspace in a Clustered Environment: (i) visualizes the environment. (ii) and (iii) show the BP and BB mode workspaces, represented by scatter plots from a side view. Scatter plots are based on 10,000 uniformly sampled joint configurations, excluding collisions.
IV-A1 PP mode
This is a trivial motion mode where the end-effector can move only up and down linearly in the task space. The IK solution lies in this mode only when $p_{t}$ is right above the modular manipulator’s base.
IV-A2 BB mode
When both units are operating bending modes, the modular manipulator can be equivalently viewed as four cylindrical links connected serially with revolute joints. There are thorough studies and several commercially ready software packages for IK of this type of structure, such as the MATLAB robotic system toolbox etc. Therefore, we also skip the detailed discussions of this mode here.
IV-A3 PB mode
In this mode, the Cartesian position of the end-effector can be mapped from the joint poses.
$\begin{cases}x=l_{2}\sin(\theta_{2})\cos(\phi_{1}+\phi_{2})\\ y=l_{2}\sin(\theta_{2})\sin(\phi_{1}+\phi_{2})\\ z=t_{1}+t_{2}+2l_{1}\cos(\frac{\theta_{1}}{2})+l_{2}\cos(\theta_{2})\end{cases}$ (1)
The result is straightforward with arithmetic.
$\begin{cases}\phi_{1}+\phi_{2}=\arctan 2(y,x)\\ \theta_{1}=2\arccos(\frac{z-t_{1}-t_{2}-\sqrt{l_{2}^{2}-x^{2}-y^{2}}}{2l_{1}})\\ \theta_{2}=\arcsin(\frac{\sqrt{x^{2}+y^{2}}}{l_{2}})\end{cases}$ (2)
Since the first unit is operating linear motion, $\theta_{1}\geq 0$, so there is only one possible choice of $\theta_{1}$ and two possible choices of $\theta_{2}$. Combining with the redundancy of $\phi_{1}$ and $\phi_{2}$, this gives the freedom to choose collision-free joint configurations. An intuitive and efficient way is iterating through discrete values of $\phi_{1}$ and the two values of $\theta_{2}$ to find a configuration with the furthest distance from obstacles $\{\mathcal{O}_{m}\}_{m=1}^{M}$. The redundancies in $\phi_{1}$, $\phi_{2}$ and $\theta_{2}$ are sometimes critical when avoiding collisions.
IV-A4 BP mode
In this mode, the Cartesian position of the end-effector can be mapped from the joint poses.
$\begin{cases}x=(l_{1}+t_{2}+2l_{2}\cos(\frac{\theta_{2}}{2}))\sin(\theta_{1})\cos(\phi_{1})\\ y=(l_{1}+t_{2}+2l_{2}\cos(\frac{\theta_{2}}{2}))\sin(\theta_{1})\sin(\phi_{1})\\ z=t_{1}+l_{1}+(l_{1}+t_{2}+2l_{2}\cos(\frac{\theta_{2}}{2}))\cos(\theta_{1})\end{cases}$ (3)
The result is also straightforward with arithmetic
$\begin{cases}\phi_{1}=\arctan 2(y,x)\\ \theta_{1}=\arcsin(\frac{\sqrt{x^{2}+y^{2}}}{\sqrt{x^{2}+y^{2}+(z-t_{1}-l_{1})^{2}}})\\ \theta_{2}=2\arccos(\frac{\sqrt{x^{2}+y^{2}+(z-t_{1}-l_{1})^{2}}-l_{1}-t_{2}}{2l_{2}})\end{cases}$ (4)
Similarly, there is also only one possible choice of $\theta_{2}$ and two possible choices of $\theta_{1}$. $\phi_{2}$ is not participated in the forward kinematics (3), which, analogous to the LR mode, leaves freedom to select collision-free joint configurations.
Although a solution for two-unit systems is provided here, efficient IK methods for modular manipulators with more units remains a challenge due to the curse of dimensionality. Further research is needed to address this issue.
Figure 6: Demonstration of Single-Joint PBT Wrist and Two-Joint Modular Manipulator (A) PBT Wrist Demonstration: A pick-and-place process is performed using the PBT wrist integrated on a UR5 robotic arm. The sequence involves the PBT wrist reaching into a large box to pick out a smaller box and stacking it in a narrow space. Subsequently, the large box is picked, reoriented, and further utilized for additional box-picking tasks. Finally, the picked box is stacked in the narrow space. (B) Modular Arm Demonstration: The two-joint PBT modular manipulator showcases its foldable and reconfigurable capabilities. The arm unfolds from within a box, reaches a specified pose, grasps a plastic bottle, and then folds back into the box.
IV-B Reachability and Manipulability Analysis
This part analyzes the reachability and manipulability of PBT joints, both as an attached tool of a collaborative robot arm and a two-unit modular manipulator [27].
Fig. 5A shows the reachable set of a single PBT joint, from which we can see that the P mode has a reachable set of a line and the B mode has sphere reachable set. Fig. 5B demonstrates the capability of PBT joints to extend a robot’s workspace in a compact environment. The PBT joint allows manipulation in a cavity area whose opening is opposite to the robot’s base. The robot end-effector is unable to enter the cavity area without the assistance of a PBT joint.
Fig. 5C visualizes the translational manipulability of distributed in the end effector’s reachable positions. Translational manipulability describes the required effort to drive the robot towards all possible task-space directions at a given configuration. It is quantified here following the definition in [28]
$\sigma(J)=\sqrt{\det(J(q)J^{T}(q))}=\sqrt{s_{1}s_{2}\dots s_{n}}$ (5)
where $J(q)$ is the translational Jacobian at configuration $q$, and $s_{1}\dots s_{n}$ are eigenvalues of matrix $J(q)J^{T}(q)$. Under PP mode, the modular manipulator can only move linearly up and down, so its reachable set is a line with zero manipulability. BP and BB modes has similar reachable set and manipulability distribution in obstacle-free environment, and has a larger volume in comparison to the workspace of PP and PB modes. In general, mode transformation makes only a minor improvement when moving in a free space.
The advantages of mode transformation becomes evident in the presence of obstacles. Prismatic mode allows PBT joints to shrink its dimensions, enhancing flexibility in compact environments, while bending mode enables the end-effector to navigate around obstacles, offering greater dexterity in clustered scenarios. Since no general quantitative measure for collision-free dexterity exists, a qualitative analysis is provided.
Figure 5D illustrates performance in a narrow tunnel. The BB mode results in a split, disconnected workspace, limiting the end-effector’s ability to transition between zones. In contrast, the BP mode creates a connected workspace, demonstrating greater flexibility in confined spaces. Figure 5E highlights performance in clustered environments. The BP mode cannot reach beyond an obstacle, whereas the BB mode achieves a significant collision-free workspace on the other side, showcasing the bending mode’s dexterity for navigating obstacles in cluttered scenarios.
V Experimental Validation
In the experimental section, we mainly demonstrate two use cases based on the PBT joints. One is the PBT wrist, and the other is the modular manipulator. The configurations of their experimental setups and the explanations and discussions of the demonstrations are included in this section.
V-A Demonstration of The One Module PBT Wrist
The PBT wrist is a small-sized PBT joint (Fig. 3). It also uses the same link configuration (Fig. 4E) to avoid singularities in force during lifting and lowering motions, with its parameters as shown in Tab. I. Unlike the previously shown PBT joint, the wrist’s revolute joint is not driven by a synchronous belt but rather by directly mounting a Dynamixel servo motor (M226) on the revolute axis, which directly drives the revolute motion [29]. Thus, the reduction ratio of the revolute joint is 1. Another identical Dynamixel servo motor is installed on the orientation-maintaining mechanism, driving the wrist’s linear motion through a gear structure with a reduction ratio of 2. Both motors are connected to the PC via a servo driver board, enabling simple control (Fig. 1B and 1C) [30].
To showcase the dexterity of the PBT wrist, we attached a suction cup at the wrist’s end to achieve pick-and-place operations in confined spaces (Fig. 6A). When vertical space is limited, the foldable and bendable structure at the end of the wrist allows effective grasping, manipulation, and placement while also increasing the variety of grasping directions.
V-B Demonstration of Two Module Manipulator
The structure of the modular manipulator is shown in Fig. 4. It uses two types of motors: the x6-40 and x4-24 motors (MyActuator Limited, Suzhou). A pair of x6-40 motors are used in the large PBT joint to drive prismatic motion, while the revolute motion of the large PBT joint is driven by a pair of x4-24 motors. Additionally, the two joints of the medium PBT joint are each driven by a single motor, with the prismatic joint and revolute joint both driven by x4-24 motors. Lastly, the revolute motors at the base of each joint are x6-40 motors. These motors are internally equipped with motor drivers, allowing direct communication with the PC for simple control of the motor groups, including synchronized control (simultaneous operation) of the motor pairs.
The motion of the module manipulator is presented in Fig. 6B. The large PBT joint provides three degrees of freedom: bottom rotation and prismatic-bending dual-modal motion. Electromagnetic clutches manage the revolute joints, disengaging during prismatic motion for passive following and engaging for revolute motion when aligned. Demonstrations showcase telescopic and grasping motions within a box, emphasizing the manipulator’s high telescoping ratio and dexterity in confined space.
VI Conclusion and Future Work
In this paper, we introduced the Prismatic-Bending Transformable (PBT) Joint, a novel robotic joint designed to enhance adaptability, dexterity, and manipulability in modular robotic systems for dynamic and unstructured environments. The PBT Joint seamlessly integrates multi-directional bending, rotation, and elongation/contraction capabilities, supported by a direction-maintenance mechanism to ensure stability and proper configuration. We demonstrated two primary applications: a single-joint PBT wrist extension for existing manipulators, and a two-joint PBT modular manipulator capable of foldable and reconfigurable configurations. Experimental results validated the PBT Joint’s ability to improve workspace coverage, enhance dexterity in confined and clustered environments, and achieve enhanced obstacle avoidance.
Future work on the PBT joint and its modular manipulator includes several promising directions to advance their capabilities. Configurations with multiple interconnected modules could address more complex tasks, enhancing the versatility of modular manipulators. Miniaturizing the PBT joint to a centimeter scale or smaller would enable fine manipulations, such as finger-level tasks or micro-robotics. Integrating PBT-based systems into advanced platforms, including humanoid robots and autonomous manipulators, could significantly expand their operational scope. Additionally, improving the design to support higher bending payloads—currently limited compared to linear payloads—and enhancing energy efficiency would broaden their industrial applications, unlocking new potential.